# Enaction for QBists

### v1

@@ 2 The Enactive Approach to Cognition
The circular organization of the sensorimotor loop—with sensory signals shaping action and action shaping sensory signals—is key to the enactive view of cognition and perception. By modulating those loops as they go around, an organism is at once behaving more or less intelligently and perceiving more or less successfully depending on how changes in those loops support or hinder autopoiesis. A hungry bacterium successfully navigating to a region with a higher sugar density can be said to be behaving intelligently and accurately perceiving the sugar gradient in its environment without ever having to form an internal model of its world. The same could be said for a human being, whose central nervous system coordinates its thirty trillion or so cells such that the body as a whole can enact its own sensorimotor loops, which both individuate and sustain the body as a single, autopoietic unit. [^3] There’s some debate about whether the term “autopoiesis” really applies to a multicellular organism like a human being, where there’s not as clear an analogue of the cell membrane, though there are other self-individuating boundaries (see footnote 2). Some enactivists prefer to describe multicellular organisms as “autonomous” rather than “autopoietic.” I’m choosing to stick with “autopoietic” here to emphasize the core enactivist ideas and the ways in which they bootstrap complex forms of mindedness from living itself. In the representational view, information from the senses is corralled into an inner picture of the world to be apprehended by the brain; in the enactive view, the brain participates in the enaction and modulation of sensorimotor loops that allow the body to interact with the environment in such a way as to remain alive.
@@
The result of perception is not an image or movie in the head, but a sensorimotor coordination, a skillful, embodied, adaptive adjustment to the world that an organism makes in order to maintain its identity and autonomy in the face of the entropy. [^5] In the enactivist view, agents are self-producing (autopoietic) systems whose existence is at all times precarious and unfinished—they must act on the world to resist entropy, which is always threatening to undo them, creating themselves again and again in the process. It’s how it draws and re-draws the subject-object divide. The agent’s perception—being active—reshapes the world and the agent, in turn, must reshape himself to adjust to that new world, which reshapes it again, and so on. The “accuracy” of a perception, in this view, lies not in some (impossible to attain) correspondence between an inner representation and an outer reality but in the resulting success or failure of the body’s coupling with the world, with all of its autopoietic ramifications. Perception, in the enactive approach, is normative.
@@ 3 The Body
Building on an epistemology of knowing how, enactivists are now adding to Ryle’s duo a third form of knowledge, one in which the enaction of the subject-object divide becomes the most ambiguous and therefore, perhaps, the most consequential. The enactivist Hanne De Jaegher calls it an engaged (or engaging) epistemology; we might call it knowing with. It requires a radical shift in perspective.
@@ 5 First, Third…Second?
In philosophy, the “I”—Descartes’ cogito—is assumed to have some privileged, private, interior access to itself; it knows itself in a way that is fundamentally different from the way it knows the world or other people. Cognitivists call this introspection. If introspection is possible, then to say “I know that the apple is red” is equivalent to saying, “I know that I know that the apple is red.” For the cognitivist, adding “I know” adds zero content, because our thoughts, beliefs, perceptions and other internal states are, in the first-person perspective, simply given to us in an immediate and transparent way. We simply have thoughts, beliefs and perceptions—we don’t have to do anything to learn out about them. We don’t have to measure them. To put it another way, a true first-person statement does not enact a subject-object divide. All collapses onto the side of the subject and its self-knowing. It’s the mirror-image of a third person statement, which is all object, no subject. The “I” is as much a hidden variable as the “it.”
@@
For enaction, contra Descartes, the “I” is not the starting point for life or for mind, let alone for philosophy or for science. Enaction is a theory of the ongoing creation of the self through the sensorimotor loops that sustain autopoiesis. But consider an infant. As babies, we can’t feed ourselves, regulate our body temperature, direct our own attention, move our eyes in synchrony, lift our heads, burp. We are not yet fully autonomous sensorimotor agents. Our caregivers have to act in coordination with us to achieve these necessary feats—which is to say they form part of the sensorimotor loops that enable our continued existence. And yet those sensorimotor loops literally constitute our selves. Our minds. The Finnish philosopher Pentti Määttänen writes [31, p.14], “Mind is not a property of the brain or even the body. Mind is a property of organism environment interaction as characterized with the loop of perception and action. Any attempt to separate one element from this loop has the consequence that mentality is lost away. Drop the loop, lose the mental.” If others form part of that loop, they form part of our minds.
Lev Vygotsky [32, p.56] famously gave the example of an infant grabbing for a toy that is just out of reach. The caregiver, seeing the baby’s intention, grabs the toy for him and hands it to him, completing the sensorimotor act. The failed reach has given birth to pointing—the baby learns that the gesture directs the caregiver’s attention in a very particular way. With gesture—be it pointing or vocalizing—the child draws the caregiver into the very sensorimotor loops that define the child as a self. Eventually the child can turn those gestures on himself and use them to regulate his own attention and coordinate his own behavior—say by talking to himself, out loud at first and eventually silently, in a process we call “thinking.”
@@ 6 The Problem of Multiple Agents
Two adults conversing with well-worn language must still enact new meanings in the course of their interaction if it’s to be a genuine conversation (and not, say, the rehearsed play-acting of small talk). One person speaks; the other responds, “Oh, you mean…?” “Not exactly,” says the first, then reformulates his words, altering, ever slightly, the original meaning. “Ah, so it’s like…” the other concedes, only in doing so he’s tweaked the meaning again. “I suppose it is,” replies the first. “It’s as if….” By the end of the conversation, a new meaning has arisen, a novel outcome that belongs to the both of them. Unlike the representational view, where communication entails the successful passage of a message from one individual to another, enactive languaging is about the negotiation of new meanings in interaction.
@@ 7 QBism and Enaction
As it happens, though, QBism does have more to say on agency, and again it intersects intriguingly with enaction. When, for instance, contrasting themselves with proponents of Relational Quantum Mechanics (RQM), QBists point out that for RQM, any physical system can be an observer—an electron, a detector, a chair—which is not the case for QBism. What, then, is QBism’s criteria for agency? Fuchs offers this [48, p.6]: “An agent is an entity that can freely take actions on parts of the world external to itself and for which the consequences of its actions matter to it.
@@
Given this view of agents, consider the problem of intersubjectivity in quantum mechanics. Without a deterministic block universe to provide a shared reality from the start, how can participators ever come to agree on the facts of the world? “In QBism,” writes Rüdiger Schack [50], “intersubjective agreement is not an automatic consequence of the quantum formalism, but a goal that agents might strive for.” It’s a statement that resonates in enactivism: “Agreements are not givens […] They must be achieved” [29, p.182]. In enactivism, they are achieved through participatory sense-making. Can this be useful to the QBist? In participatory sense-making, the interaction between agents has its own autonomy that co-exists along with the autonomies of the individual agents. In other words, the same process by which the individual forges his own identity takes place at the level of the interaction. Could something similar occur between agents in quantum mechanics?
@@
Whether this move would get us somewhere is hard to say, but the point is, it’s only in an enactivist framework that we can even pose such a suggestion. For if we took agents to be forming metaphysically private internal models of the world, then two agents interacting with one another couldn’t possibly produce a “third agent” between them. Genuine intersubjectivity would be impossible from the start.
Conversations between enactivists and QBists are just beginning [51], and it’s impossible to say where they may lead. But according to both research programs, their interaction is bound to produce something novel. If QBism is ultimately a theory of how the world is created and recreated through interactions with an agent, enactivism is a theory of how the agent is created and recreated through interactions with the world. One can’t help but feel that each is half of the same participatory story.
@@ 8 Conclusion: Toward a Unified Approach
QBism and enactivism are theories of how Wheeler’s chalk line—the subject-object divide, which was, in classical physics, pre-given and absolute—is enacted and renegotiated interaction by interaction. Both theories reject representationalism, which requires that fixed divide. (“We propose as a name the term enactive to emphasize the growing conviction that cognition is not the representation of a pregiven world by a pregiven mind” [7]. “One of the main conclusions of QBism is that quantum mechanics is not representational” [50].) For QBism, quantum measurement is the enactment of that boundary; for enactivism, cognition is the enactment of that boundary; both are processes that are forever unfinished. For the former, the world is created on the fly; for the latter, the mind is created on the fly. QBism chipped away at the classical object, while enactivism chipped away at the classical subject, and between them they opened a new place in the middle—in the drawing of the divide itself—for them to meet.
Today, both theories are still seeking their ontologies, and it’s for that, perhaps, that they need one another most. QBism accounts for the participatory becoming of the world, but often leaves agents as primitives in the theory; enactivism accounts for the participatory becoming of agents, but often leaves the world as a primitive. Just as QBism would fall into incoherence if it understood agency, belief, experience, and normativity in representational, cognitivist terms, enactivism would fall into incoherence if it understood agents to be acting within a world governed by classical, non-participatory physics. Enactivism needs the QBist understanding of the world as crucially as QBism needs the enactivist understanding of the agent. (Perhaps, then, these notes should be followed by a companion set: QBism for Enactivists.)
“The classical representationalist picture of thought and how the mind works is so powerful and so deeply rooted in our self-understanding that it is hardly likely to ever be dislodged,” Johnson writes [27, p.132]. The same, surely, can be said of the classical realist picture of matter and how the universe works. QBism and enactivism both stand as challenges to our most deeply entrenched notions of world and self, the very foundations of reality. They face uphill battles, to say the least. But the point is, it’s the same battle. Both are fighting against the same 17th century Cartesian split, the same dichotomy of subject and object. Trying to convince others of QBism or enactivism in isolation may be tough going precisely because it leaves half the story untold. Together, though, QBism and enactivism make for powerful allies, as they can offer, for the first time, a unified post-Cartesian metaphysics, one that finally brings self and world back together again.

### v2

@@ 2 The Enactive Approach to Cognition
The circular organization of the sensorimotor loop—with sensory signals shaping action and action shaping sensory signals—is key to the enactive view of cognition and perception. By modulating those loops as they go around, an organism is at once behaving more or less intelligently and perceiving more or less successfully depending on how changes in those loops support or hinder autopoiesis. A hungry bacterium successfully navigating to a region with a higher sugar density can be said to be behaving intelligently and accurately perceiving the sugar gradient in its environment without ever having to form an internal model of its world. The same could be said for a human being, whose central nervous system coordinates its thirty trillion or so cells such that the body as a whole can enact its own sensorimotor loops, which both individuate and sustain the body as a single, autopoietic unit. [^3] There’s some debate about whether the term “autopoiesis” really applies to a multicellular organism like a human being, where there’s not as clear an analogue of the cell membrane, though there are other self-individuating boundaries (see footnote 2). Some enactivists prefer to describe multicellular organisms as “autonomous” rather than “autopoietic.” I’m choosing to stick with “autopoietic” here to emphasize the core enactive ideas and the ways in which they bootstrap complex forms of mindedness from living itself. In the representational view, information from the senses is corralled into an inner picture of the world to be apprehended by the brain; in the enactive view, the brain participates in the enaction and modulation of sensorimotor loops that allow the body to interact with the environment in such a way as to remain alive.
@@
The result of perception is not an image or movie in the head, but a sensorimotor coordination, a skillful, embodied, adaptive adjustment to the world that an organism makes in order to maintain its identity and autonomy in the face of the entropy. [^5] In the enactive view, agents are self-producing (autopoietic) systems whose existence is at all times precarious and unfinished—they must act on the world to resist entropy, which is always threatening to undo them, creating themselves again and again in the process. It’s how it draws and re-draws the subject-object divide. The agent’s perception—being active—reshapes the world and the agent, in turn, must reshape himself to adjust to that new world, which reshapes it again, and so on. The “accuracy” of a perception, in this view, lies not in some (impossible to attain) correspondence between an inner representation and an outer reality but in the resulting success or failure of the body’s coupling with the world, with all of its autopoietic ramifications. Perception, in the enactive approach, is normative.
@@ 3 The Body
Building on an epistemology of knowing how, enactivists are now adding to Ryle’s duo a third form of knowledge, one in which the enaction of the subject-object divide becomes the most ambiguous and therefore, perhaps, the most consequential. The enactivist Hanne De Jaegher calls it an engaged (or engaging) epistemology; we might call it knowing with. And it requires a radical shift in perspective.
@@ 5 First, Third…Second?
In philosophy, the “I”—Descartes’ cogito—is assumed to have some privileged, private, interior access to itself; it knows itself in a way that is fundamentally different from the way it knows the world or other people. Cognitivists call this introspection. If introspection is possible, then to say “I know that the apple is red” is equivalent to saying, “I know that I know that the apple is red.” For the cognitivist, adding “I know” adds zero content, because our thoughts, beliefs, perceptions and other internal states are, in the first-person perspective, simply given to us in an immediate and transparent way. We simply have thoughts, beliefs and perceptions—we don’t have to do anything to learn about them. We don’t have to measure them. To put it another way, a true first-person statement does not enact a subject-object divide. All collapses onto the side of the subject and its self-knowing. It’s the mirror-image of a third person statement, which is all object, no subject. The “I” is as much a hidden variable as the “it.”
@@
For enaction, contra Descartes, the “I” is not the starting point for life or for mind, let alone for philosophy or for science. Enaction is a theory of the ongoing creation of the self through the sensorimotor loops that sustain autopoiesis. But consider an infant. As babies, we can’t feed ourselves, regulate our body temperature, direct our own attention, move our eyes in synchrony, lift our heads, burp. We are not yet fully autonomous sensorimotor agents. Our caregivers have to act in coordination with us to achieve these necessary feats—which is to say, they form part of the sensorimotor loops that enable our continued existence. And yet those sensorimotor loops literally constitute our selves. Our minds. The Finnish philosopher Pentti Määttänen writes [31, p.14], “Mind is not a property of the brain or even the body. Mind is a property of organism environment interaction as characterized with the loop of perception and action. Any attempt to separate one element from this loop has the consequence that mentality is lost away. Drop the loop, lose the mental.” If others form part of that loop, they form part of our minds.
Lev Vygotsky [32, p.56] famously gave the example of an infant grabbing for a toy that is just out of reach. The caregiver, seeing the baby’s intention, grabs the toy and hands it to him, completing the sensorimotor act. The failed reach has given birth to pointing—the baby learns that the gesture directs the caregiver’s attention in a very particular way. With gesture—be it pointing or vocalizing—the child draws the caregiver into the very sensorimotor loops that define the child as a self. Eventually the child can turn those gestures on himself and use them to regulate his own attention and coordinate his own behavior—say, by talking to himself, out loud at first and eventually silently, in a process we call “thinking.”
@@ 6 The Problem of Multiple Agents
Two adults conversing with well-worn language must still enact new meanings in the course of their interaction if it’s to be a genuine conversation (and not, say, the rehearsed play-acting of small talk). One person speaks; the other responds, “Oh, you mean…?” “Not exactly,” says the first, then reformulates his words, altering, ever so slightly, the original meaning. “Ah, so it’s like…” the other concedes, only in doing so he’s tweaked the meaning again. “I suppose it is,” replies the first. “It’s as if….” By the end of the conversation, a new meaning has arisen, a novel outcome that belongs to the both of them. Unlike the representational view, where communication entails the successful passage of a message from one individual to another, enactive languaging is about the negotiation of new meanings in interaction.
@@ 7 QBism and Enaction
As it happens, though, QBism does have more to say on agency, and again it intersects intriguingly with enaction. When, for instance, contrasting themselves with proponents of Relational Quantum Mechanics (RQM), QBists point out that for RQM, any physical system can be an observer—an electron, a detector, a chair—which is not the case for QBism. What, then, is QBism’s criteria for agency? Fuchs offers this [48, p.6]: “An agent is an entity that can freely take actions on parts of the world external to itself and for which the consequences of its actions matter to it.”
@@
Given this view of agents, consider the problem of intersubjectivity in quantum mechanics. Without a deterministic block universe to provide a shared reality from the start, how can participators ever come to agree on the facts of the world? “In QBism,” writes Rüdiger Schack [50], “intersubjective agreement is not an automatic consequence of the quantum formalism, but a goal that agents might strive for.” It’s a statement that resonates with the enactive approach: “Agreements are not givens […] They must be achieved” [29, p.182]. In enaction, they are achieved through participatory sense-making. Can this be useful to the QBist? In participatory sense-making, the interaction between agents has its own autonomy that co-exists along with the autonomies of the individual agents. In other words, the same process by which the individual forges his own identity takes place at the level of the interaction. Could something similar occur between agents in quantum mechanics?
@@
Whether this move would get us somewhere is hard to say, but the point is, it’s only in an enactive framework that we can even pose such a suggestion. For if we took agents to be forming metaphysically private internal models of the world, then two agents interacting with one another couldn’t possibly produce a “third agent” between them. Genuine intersubjectivity would be impossible from the start.
Conversations between enactivists and QBists are just beginning [51], and it’s impossible to say where they may lead. But according to both research programs, their interaction is bound to produce something novel. If QBism is ultimately a theory of how the world is created and recreated through interactions with an agent, enaction is a theory of how the agent is created and recreated through interactions with the world. One can’t help but feel that each is half of the same participatory story.
@@ 8 Conclusion: Toward a Unified Approach
QBism and enaction are theories of how Wheeler’s chalk line—the subject-object divide, which was, in classical physics, pregiven and absolute—is enacted and renegotiated interaction by interaction. Both theories reject representationalism, which requires that fixed divide. (“We propose as a name the term enactive to emphasize the growing conviction that cognition is not the representation of a pregiven world by a pregiven mind” [7]. “One of the main conclusions of QBism is that quantum mechanics is not representational” [50].) For QBism, quantum measurement is the enactment of that boundary; for enaction, cognition is the enactment of that boundary; both are processes that are forever unfinished. For the former, the world is created on the fly; for the latter, the mind is created on the fly. QBism chipped away at the classical object, while enaction chipped away at the classical subject, and between them they opened a new place in the middle—in the drawing of the divide itself—for them to meet.
Today, both theories are still seeking their ontologies, and it’s for that, perhaps, that they need one another most. QBism accounts for the participatory becoming of the world, but often leaves agents as primitives in the theory; enaction accounts for the participatory becoming of agents, but often leaves the world as a primitive. Just as QBism would fall into incoherence if it understood agency, belief, experience, and normativity in representational, cognitivist terms, enaction would fall into incoherence if it understood agents to be acting within a world governed by classical, non-participatory physics. Enaction needs the QBist understanding of the world as crucially as QBism needs the enactive understanding of the agent. (Perhaps, then, these notes should be followed by a companion set: QBism for Enactivists.)
“The classical representationalist picture of thought and how the mind works is so powerful and so deeply rooted in our self-understanding that it is hardly likely to ever be dislodged,” Johnson writes [27, p.132]. The same, surely, can be said of the classical realist picture of matter and how the universe works. QBism and enaction both stand as challenges to our most deeply entrenched notions of world and self, the very foundations of reality. They face uphill battles, to say the least. But the point is, it’s the same battle. Both are fighting against the same 17th century Cartesian split, the same dichotomy of subject and object. Trying to convince others of QBism or the enactive approach in isolation may be tough going precisely because it leaves half the story untold. Together, though, QBism and enaction make for powerful allies, as they can offer, for the first time, a unified post-Cartesian metaphysics, one that finally brings self and world back together again.